# The Mixed Birth-death/death-Birth Moran Process

### v2

@@ A.2 Martingales and Drifts
Now $\phi^{\prime}(Y_{i}^{\prime})$ satisfies the conditions in Theorem˜A.2 with the same parameters, since for all $Y_{i}^{\prime}$ such that $\phi^{\prime}(Y_{i}^{\prime})\notin\{0,k_{1}\}$, the conditions holds by assumption, and if $\phi^{\prime}(Y_{i}^{\prime})=k_{1}$, then $\phi^{\prime}(Y_{i+1}^{\prime})=0$ so the conditions also hold. Applying Theorem˜A.2 gives the claim. ∎
@@
Finally, we show the FPRAS construction. The idea is just to repeatedly perform experiments and drop the experiments that took too long to terminate.
@@
If any run hits the cutoff $T$ without absorbing, abort and return an error. Otherwise output
@@
  • If all $N$ runs absorb before $T$, then $X_{1},\dots,X_{N}$ are i.i.d. $\mathrm{Bernoulli}(f)$. Hoeffding’s inequality implies
    $\mathbb{P}(|\hat{f}-f|>\varepsilon f)\leqslant 2\exp(-2\varepsilon^{2}f^{2}N)\leqslant 2\exp\left(-2\varepsilon^{2}c_{\mathrm{fp}}^{2}n^{-2c_{1}}N\right)\leqslant\tfrac{1}{8},$
@@
    By a union bound, the probability that any of the $N$ runs is cut off is at most $1/8$.
Combining the two events, with probability at least $1-1/8-1/8\geqslant 3/4$, the algorithm returns a value within a factor $1\pm\varepsilon$ of $f$. ∎
@@ Appendix B Regular Graphs
$\psi_{u,v}(S)=\lambda\left[\frac{1}{n}\cdot\frac{1}{\deg(u)}-\frac{1}{n}\cdot\frac{1}{\deg(v)}\right]+(1-\lambda)\left[\frac{1}{n}\cdot\frac{1}{\deg(v)}-\frac{1}{n}\cdot\frac{1}{\deg(u)}\right]=0.$
Thus, $\phi(S)$ is a martingale. The theorem follows from Theorem˜2.6 and 2.3.
We now bound the absorption time. Let the graph $G$ be $d$-regular. Using the exact same argument as in the proof of Theorem˜3.1, we obtain that the absorption time is $O(n^{3}d)$ when $r=1$. It remains to show the case where $r\neq 1$. We establish the result for $r>1$. The case $r<1$ is symmetric by swapping the roles of mutants and residents.
@@
For the dB component, we perform an analysis similar to that in the proof of Theorem˜4.7.
@@
Since $\phi(V)=n$, applying Theorem˜2.5 yields an absorption time of $\frac{r}{r-1}O(n^{3}d)=O_{r}(n^{4})$. ∎

### v3

@@ A.2 Martingales, Drifts, and Approximation Schemes
Now $\phi^{\prime}(Y_{i}^{\prime})$ satisfies the conditions in Theorem A.2 with the same parameters, since for all $Y_{i}^{\prime}$ such that $\phi^{\prime}(Y_{i}^{\prime})\notin\{0,k_{1}\}$, the conditions holds by assumption, and if $\phi^{\prime}(Y_{i}^{\prime})=k_{1}$, then $\phi^{\prime}(Y_{i+1}^{\prime})=0$ so the conditions also hold. Applying Theorem A.2 gives the claim. ∎
@@
Finally, we show the FPRAS construction via standard techniques. The idea is to run the experiments enough times and report the empirical mean, while aborting if any run took too long.
@@
If any run hits the cutoff $T$ without absorbing, abort and return an error $\hat{f}$$\bot$. Otherwise, output
@@
  • Consider a variant of the algorithm that always simulates until absorption with output $\tilde{f}=\frac{1}{N}\sum_{i=1}^{N}X_{i}$. Then $X_{1},\dots,X_{N}$ are i.i.d. $\mathrm{Bernoulli}(f)$. Hoeffding’s inequality implies
    $\mathbb{P}(|\tilde{f}-f|>\varepsilon f)\leqslant 2\exp(-2\varepsilon^{2}f^{2}N)\leqslant 2\exp\left(-2\varepsilon^{2}C_{\mathrm{fp}}^{2}n^{-2c_{1}}N\right)\leqslant\tfrac{1}{8},$
@@
    We have $\hat{f}=\tilde{f}$ if and only if no run is cut off. By a union bound, $\mathbb{P}(\hat{f}\neq\tilde{f})\leqslant 1/8$.
Combining the two events, the failure probability is at most
$\mathbb{P}(|\tilde{f}-f|>\varepsilon f)\leqslant\mathbb{P}(|\tilde{f}-f|>\varepsilon f)+\mathbb{P}(\hat{f}\neq\tilde{f})\leqslant 1/4.\qed$
One could amplify the success probability to $1-\delta$ by running the algorithm $O(\log(1/\delta))$ times and take the median (see, e.g. [vazirani2010approx]).
@@ Appendix B Regular Graphs
$\psi_{u,v}(S)=\lambda\left[\frac{1}{n}\cdot\frac{1}{\deg_{u}}-\frac{1}{n}\cdot\frac{1}{\deg_{v}}\right]+(1-\lambda)\left[\frac{1}{n}\cdot\frac{1}{\deg_{v}}-\frac{1}{n}\cdot\frac{1}{\deg_{u}}\right]=0.$
Thus, $\phi(S)$ is a martingale. The theorem follows from Theorem 2.6 and 2.3.
We now bound the absorption time. Let the graph $G$ be $d$-regular. Using the exact same argument as in the proof of Theorem 3.1, we obtain that the absorption time is $O(n^{3}d)$ when $r=1$. It remains to show the case where $r\neq 1$. We establish the result for $r>1$. The case $r<1$ is symmetric by swapping the roles of mutants and residents.
@@
For the dB component, we perform an analysis similar to that in the proof of Theorem 4.7.
@@
Since $\phi(V)=n$, applying Theorem 2.5 yields an absorption time of $\frac{r}{r-1}O(n^{3}d)=O_{r}(n^{4})$. ∎